If a 320. mg sample of technetium−99m is used for a diagnostic procedure, how much Tc−99m remains after each interval: (a) 6.0 h; (b) 2 days? (assume that a day is an exact value) The half-life of technetium−99m is 6.0 hours.
1
Expert's answer
2020-06-22T09:15:48-0400

Explanation:

The basic formula for half life equations is

A=P(1/2)t/h

A= amount remaining

P= original amount

t= elapsed time

h=half life time

a) P=320, t=6,h=6

A=320(1/2)6/6

=160mg

b) P=320, t=48hrs, h=6hrs

A=320(1/2)48/6

=1.25mg